OpenStudy (anonymous):

-304+246+(189)+107 show steps

OpenStudy (anonymous):

-304+246=-58 -58+189=247 247+107=354

OpenStudy (anonymous):

if you are looking for an optimization method there is something called the transfer method, basically -304+246+(189)+107 = -304+235+(200)+107 =-300+235+(200)+100 +7-4